2x+5=47-4x
2x+4x=47-5
6x=45
x=45/6

Found 2 solutions by nvegupta, lenny460:
Answer by nvegupta(2)

2x+5=47-4x
2x+4x=47-5
6x=45
x=45/6

Answer by lenny460(1073)

2x + 5 = 47 - 4x

Add 4x to each side of the equation
2x + 5 + 4x = 47 - 4x + 4x

Combine like terms:
2x + 4x = 6x
-4x + 4x = 0

6x + 5 = 47
Subtract 5 from each side of the equation
6x + 5 - 5 = 47 - 5
Combine like terms:
5 - 5 = 0
47 - 5 = 42

6x = 42
Divide each side by 6
6x/6 = 42/6
x = 7

The Answer:
x = 7

Check:

2x + 5 = 47 - 4x
2(7) + 5 = 47 - 4(7)
14 + 5 = 47 - 28
14 + 5 = 19
47 - 28 = 19
19 = 19


obuong3@aol.com